OpenStudy (anonymous):

Henry grabs one shoe at a time at random without replacement from a box of ten shoes in his closet. The box contains 2 brown shoes, 4 black shoes, and 4 sneakers. What is the probability that the first two shoes he chooses will be brown? Round to the nearest percent if necessary. 2% 20% 22% 40%

OpenStudy (wonderwomen):

20%

OpenStudy (anonymous):

are u sure:(

OpenStudy (wonderwomen):

yes i am

OpenStudy (anonymous):

\[\frac{2}{10}\cdot\frac{1}{9}\]

OpenStudy (anonymous):

ok if u say so :(

OpenStudy (wonderwomen):

that is not a even a answer sithsandgiggles

OpenStudy (anonymous):

ik right?

OpenStudy (anonymous):

Yes it is. \(\dfrac{2}{90}=0.02222222...\)

OpenStudy (anonymous):

thats not a answere.:)

OpenStudy (anonymous):

"Round to the nearest percent if necessary." Do you know what that entails? A decimal is not a percent, but you can easily convert. Multiply \(0.02\) by a 100%.

OpenStudy (wonderwomen):

but when you put it into a percent than it is 22%

OpenStudy (wonderwomen):

i mean .022%

OpenStudy (anonymous):

\(0.02\not=0.02\%\) "Percent" means "parts per 100". 0.02 is equivalent to two hundredths, or 2 parts per 100, or 2%.

OpenStudy (anonymous):

:(?

OpenStudy (anonymous):

ok?

OpenStudy (wonderwomen):

um i know how to do this i am a 7th grader

OpenStudy (anonymous):

Well don't let me waste your time if you're so sure my answer is wrong.